OpenStudy (johnnydicamillo):

Suppose that f(1) = 2, f(4) = 5, f'(4) =3 and f'' is continuous. Find the value of

OpenStudy (johnnydicamillo):

\[\int\limits_{1}^{4}xf''(x)dx\]

OpenStudy (anonymous):

Integrate by parts: \[\begin{matrix}u=x&&&dv=f''(x)\,dx\\ du=dx&&&v=f'(x)\end{matrix}\] which gives \[\int_1^4xf''(x)\,dx=\left[xf'(x)\right]_1^4-\int_1^4f'(x)\,dx=4f'(4)-f'(1)-(f(4)-f(1))\] Are you given \(f'(1)\) ?

OpenStudy (johnnydicamillo):

no f'(1) is not given

OpenStudy (johnnydicamillo):

a peer told me the professor added that f'(1) = 0

OpenStudy (johnnydicamillo):

okay then I would just need to plug it in.

OpenStudy (anonymous):

Well that's convenient. I was starting to wonder if there was some theorem that would have allowed you to determine the value using what you were given.

OpenStudy (johnnydicamillo):

Maybe there is, but thanks. How did you know to use integration by parts?

OpenStudy (anonymous):

Mainly by the fact that you're given values of the derivative at the endpoints of the interval \([1,4]\). IBP provides a way of reducing the order of the derivative and changing the integral of the second derivative to an expression containing an integral of the first derivative.

OpenStudy (anonymous):

The form of the given integral is also highly reminiscent of the integral \(\int xe^x\,dx\), which is usually evaluated with IBP.

OpenStudy (johnnydicamillo):

Ah, thanks for the explanation!

OpenStudy (anonymous):

You're welcome